OpenStudy (anonymous):

Solve 3x+2y=7 3x^2-2y^2=25

OpenStudy (anonymous):

ANSWER = {(3,-1),(11,-13)}

hartnn (hartnn):

find x from 3x+2y=7 put it in 2nd equation

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

\[x=\frac{ 7-2y }{ 3 } RIGHT?\]

OpenStudy (anonymous):

49-28y+4y^2/3 -2y^2=25

OpenStudy (anonymous):

am i doing right

hartnn (hartnn):

yes. (49-28y+4y^2)/3 -2y^2=25 first multiply 3 to eliminate denominator

OpenStudy (anonymous):

25x3=75

hartnn (hartnn):

and -2y^2 *3 = -6y^2

OpenStudy (anonymous):

2y^2-28y-26=0

hartnn (hartnn):

how u get +2y^2?? i get -2y^2

OpenStudy (anonymous):

4y^2-2y^2

hartnn (hartnn):

4y^2-6y^2

OpenStudy (anonymous):

why are u multiplying 3 by -2y^2 we are cross multiplying and it will be 25x3=75

hartnn (hartnn):

i am multiplying 3 throught , to each term

OpenStudy (anonymous):

\[\frac{ 49-28y+4y^2 }{ 3 }-2y^2=25\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[2y^2+28y+26=0\]

hartnn (hartnn):

yes.

OpenStudy (anonymous):

y=-4 and y=13

OpenStudy (anonymous):

sorry -13

hartnn (hartnn):

i m getting -1,-13

OpenStudy (anonymous):

sorry i am getting -1 and -13

OpenStudy (anonymous):

i have got the answers thank u all